OpenStudy (jackfirepower9):

I dont know what to do when I get -b = 22a When solving 7a - b = 15a????

OpenStudy (welshfella):

to get b in terms of a divide both sides by -1

OpenStudy (jackfirepower9):

ok let me try

OpenStudy (jackfirepower9):

im not typing if it says

OpenStudy (welshfella):

you have made a mistake in the first line

OpenStudy (jackfirepower9):

uh oh

OpenStudy (jackfirepower9):

aww subtract

OpenStudy (welshfella):

7a - b = 15a subtract 7a from both sides 7a - 7a - b = 15a - 7a -b = 8a now divide both sides by -1.

OpenStudy (jackfirepower9):

its B

OpenStudy (jackfirepower9):

D

OpenStudy (jackfirepower9):

is it D

OpenStudy (welshfella):

oh you want a in terms of b -b = 8a divide both sides by 8 a = -b/8

OpenStudy (jackfirepower9):

so whats the answer

OpenStudy (welshfella):

that is

OpenStudy (jackfirepower9):

that is what?

OpenStudy (welshfella):

a = -b/8 is the answer

OpenStudy (jackfirepower9):

ok thx

OpenStudy (welshfella):

yw